# From Parameters to Behaviors: Unsupervised Compression of the Policy Space

### v1

@@ Appendix B Experimental Details
Reproducibility.  In MC, we perform two main experiments. First, we study different configurations by creating 27 different datasets. We seed all the steps of the pipeline with seeds 0 through 26. In order, we use seeds 0-8 for Small policies, 9-17 for Medium policies, and 18-26 for Large policies. In each batch, the first three seeds are used for datasets of 10k policies, the next three for datasets of 50k, and the last three for datasets of 100k policies. The second experiment is focused on datasets of 10k Medium policies, and its run with seeds starting from 100. In Reacher, we focus on datasets of 100k Medium policies with seeds starting from 0.
@@
Latent Behavior Manifolds.  To complement the results in the main paper, we provide an extensive set of visualizations of the learned latent behavior manifolds. These plots illustrate how the latent representations organize policies across different tasks, policy sizes, and encoding dimensions. They cover both environments studied in this work—Mountain Car (MC) and Reacher (RC)—and show how the manifold structure emerges consistently across settings. The visualizations serve two purposes: (i) to confirm that the latent space captures meaningful behavioral structure qualitatively, and (ii) to demonstrate the consistency of this organization across seeds and settings. For clarity in the main text, we only reported a subset of representative plots; here, to enable a more thorough inspection and reproducibility, we visualize one seed per configuration in MC and all seeds for RC. Each figure shows the latent spaces for 1, 2, and 3 dimensions for all tasks. The tasks are ordered from left to right as such: standard, left, speed, height for MC, and speed, clockwise, c-clockwise, and radial for RC.

### v2

@@ Appendix B Experimental Details
Reproducibility.  In MC, we perform two main experiments. First, we study different configurations by creating 27 different datasets. We seed all the steps of the pipeline with seeds 0 through 26. In order, we use seeds 0-8 for Small policies, 9-17 for Medium policies, and 18-26 for Large policies. In each batch, the first three seeds are used for datasets of 10k policies, the next three for datasets of 50k, and the last three for datasets of 100k policies. The second experiment focuses on datasets of 10k Medium policies, and it is run with seeds starting from 100. In Reacher, we focus on datasets of 100k Medium policies with seeds starting from 0. In Hopper and HalfCheetah, we focus on datasets of 10k Medium policies with seeds starting from 0.
@@
Policy Representation Range.  We analyze the behavioral distributions of randomly sampled Small and Large policies in the MC environment (Figure 8). For simpler tasks (standard, left, and speed), both architectures cover a similar reward range, though their probability distributions differ. However, in the more complex height task, Large policies exhibit a significantly broader support, reaching higher maximum rewards than their Small counterparts. While this observation relies on random weight sampling and is limited to a single environment, it aligns with theoretical findings on the relationship between parameter count and network expressivity (Montúfar et al., 2014; Raghu et al., 2017; Bahri et al., 2024), suggesting that larger networks are naturally capable of representing a richer diversity of behaviors.
Figure 6: Reward distribution comparison of datasets of 10k Small (blue) and Large (orange) policies in MC. The $y$-axis uses a logarithmic scale. We report the average and 95% CI over 5 seeds.
Latent Behavior Manifolds.  To complement the results in the main paper, we provide an extensive set of visualizations of the learned latent behavior manifolds. These plots illustrate how the latent representations organize policies across different tasks, policy sizes, and encoding dimensions. They cover both environments studied in this work—Mountain Car (MC) and Reacher (RC)—and show how the manifold structure emerges consistently across settings. The visualizations serve two purposes: (i) to confirm that the latent space captures meaningful behavioral structure qualitatively, and (ii) to demonstrate the consistency of this organization across seeds and settings. For clarity in the main text, we only reported a subset of representative plots; here, to enable a more thorough inspection and reproducibility, we visualize one seed per configuration in MC and all seeds for RC. Each figure shows the latent spaces for 1, 2, and 3 dimensions for all tasks. The tasks are ordered from left to right as follows: standard, left, speed, height for MC, and speed, clockwise, c-clockwise, and radial for RC.